001 10.0 points A car accelerates uniformly from rest and reaches a speed of 14.9 m/s in 11.5 s. The diameter of a tire is 71.3 cm. Find the number of revolutions the tire makes during this motion, assuming no slip- ping. Answer in units of rev.
Answer #1

v² - (vo)² = 2ax

(14.9 m/s)² - 0 = 2(14.9 m/s / 11.5 s)(x)

x = 85.675 m

n = (85.675 m) / (π(0.713 m))

n = 38.25 rev